The Smith Brothers ages multiplied together give the answer 2184. Tom is 1 year older than john. Michael is 1 year older than Tom. How old are they?
Answer by swincher4391(1107)

Let John's age be j. Tom be t And Michael be m.
j+1 = t Since Tom is 1 year older than John
t+1 = m Since Michael is 1 year older than Tom.
Solve for t in the 2nd to get t = m-1
Then j+1 = m-1
So j+2 = m
So we have j, j+1, j+2
Then j(j+1)(j+2) = 2184
j^3 + 3j^2 + 2j = 2184
j^3 + 3j^2 + 2j -2184 = 0
We get two complex answers and one real one... that being j = 12.
So John is 12.
Tom is 13.
Michael is 14.
Check:
12*13*14 = 2184 Yes.